OpenStudy (lacedupcountry):

the sum of three consecutive integers is 267, what is the largest integer? i don't understand how to do it.

OpenStudy (wolf1728):

okay let's word it differently (basically a new problem) Three consecutive integers sum to 9. What is the largest integer?

OpenStudy (lacedupcountry):

3?

OpenStudy (wolf1728):

No it's four but at least you understand the problem right?

OpenStudy (wolf1728):

2 + 3 + 4 = 9

OpenStudy (lacedupcountry):

oh okay i feel super dumb now, i get it haha thanks

OpenStudy (lacedupcountry):

so i can do 100+100+67 ?

OpenStudy (wolf1728):

okay so now we have the problem where they add to 267 so we'll call the first integer n, the next one is (n+1) and the final one is (n+2) I think you can see that right?

OpenStudy (wolf1728):

no those are not consecutive

OpenStudy (lacedupcountry):

so the third integer would be (n+3) ?

OpenStudy (wolf1728):

actually n+2 since the first one is n.

OpenStudy (lacedupcountry):

Okay, now what?

OpenStudy (wolf1728):

Okay maybe we should work on the problem where they add up to 9. First integer is n, then n+1 then n+2 we have the equation n + n+1 + n +2 = 9

OpenStudy (lacedupcountry):

i get it now thanks

OpenStudy (wolf1728):

so you now have the equation where n + n+1 + n+2 = 267 right?

OpenStudy (lacedupcountry):

n=91?

OpenStudy (wolf1728):

No but very close

OpenStudy (lacedupcountry):

Sorry hang on i messed up in my work lol

OpenStudy (lacedupcountry):

n=88

OpenStudy (wolf1728):

Very good !! :-)

OpenStudy (lacedupcountry):

yay haha Thankyou! :)

OpenStudy (wolf1728):

And to answer the question the largest integer is 90.

OpenStudy (lacedupcountry):

Got it

OpenStudy (wolf1728):

okay !!! glad to help you out